A farmer bought 100 animals for $100. Sheep were $10 each, pigs were $2
each, and chickens were $.50 each. How many of each did she buy?

Answer by Edwin McCravy(20054)

A farmer bought 100 animals for $100. Sheep were $10 each, pigs were $2
each, and chickens were $.50 each. How many of each did she buy?
  S +  P +   H = 100
10S + 2P + .5H = 100

TWO SOLUTIONS:

  1 sheep  1×$10 =  $10
 27 pigs  27×$2  =  $54
 72 hens  .5×72  =  $36
----------------------
100 animals      = $100

  4 sheep  4×$10 =  $40
  8 pigs   8×$2  =  $16
 88 hens  .5×88  =  $44
----------------------
100 animals      = $100   


Edwin